# JASA-supplement

### v3

@@ 1 Introduction
Current computational approaches for MVN probabilities are mostly based on the separation of variable (SOV) method (Genz, 1992). SOV applies a series of transformations such that the MVN probability can be written as an expectation of quantities with low variance, allowing an approximation of the expectation as the average of Monte Carlo (MC) samples. SOV consists of two stages: in the pre-processing stage, the Cholesky factor of the $n\times n$ MVN covariance matrix is computed in $\mathcal{O}(n^{3})$ time; in the MC sampling stage, $N$ samples are drawn, each in $\mathcal{O}(n^{2})$ time. The resulting SOV time complexity is $\mathcal{O}(n^{3}+Nn^{2})$. To improve upon the original SOV method, Gibson et al. (1994) introduced a univariate variable reordering technique used in the pre-processing stage that produces the Cholesky factor and further reduces the MC variance. Based on the series of integrand transformations used in the SOV method, Botev (2017) used importance sampling with minimax exponentially tilted (MET) normal proposal densities during the MC sampling stage to reduce variance. This results in substantially higher accuracy for tail MVN probabilities than SOV, but it requires iterative optimization of the proposal density in the pre-processing stage at a per-iteration complexity of $\mathcal{O}(n^{3})$; in practice, this optimization can be more expensive than the MC sampling for moderately large $n$ (e.g., $n>1{,}000$). Sampling from truncated MVN (TMVN) distributions is closely connected to approximating MVN probabilities (Botev, 2017). Specifically, the integrand in MET (or SOV) can be decomposed into two parts that can be viewed as a proposal density and a density ratio, which can be used for importance sampling from TMVN distributions at the same cost as estimating the corresponding MVN probability. MET is currently the state-of-the-art in terms of achieving the highest accuracy for estimating MVN probabilities and the highest acceptance rate for sampling TMVN distributions.
@@ 2 Importance-Sampling Views of SOV and MET
[…]
@@
[…]
@@
where $\bm{\gamma}$ are the parameters for the proposal density used in MET, and $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function (PDF) and the cumulative probability function (CDF) of the univariate standard normal distribution, respectively. From the perspective of importance sampling, (3) and (5) correspond to the SOV and the MET methods, respectively. Specifically, $q^{B}(\mathbf{y};\bm{\gamma})$ is the proposal density and $h^{B}(\mathbf{y};\bm{\gamma})$ is the density ratio between $\phi_{n}(\mathbf{y};\mathbf{I}_{n})$ truncated in $\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}$ and $q^{B}(\mathbf{y};\bm{\gamma})$, up to a normalizing constant. MET samples $\mathbf{y}$ from $q^{B}(\mathbf{y};\bm{\gamma})$ and then approximates $E_{q^{B}(\mathbf{y};\bm{\gamma})}[h^{B}(\mathbf{y};\bm{\gamma})]$ by the sample average. Unlike TMVN distributions, $q^{B}(\mathbf{y};\bm{\gamma})$ can be sampled as a sequence of $n$ univariate truncated normal distributions as implied by (8). Hence, we can apply accept-reject sampling to TMVN distributions by using $q^{B}(\mathbf{y};\bm{\gamma})$ as the proposal density to draw samples from $\phi_{n}(\mathbf{x};\bm{\Sigma})$ truncated in $[\mathbf{a},\mathbf{b}]$. SOV is a special case of MET when the exponential tilting parameters $\bm{\gamma}=\mathbf{0}$. Using
@@ 3.1 Conditional mean and variance under Vecchia
$\eta_{i}$ :$\vcentcolon=E[x_{i}|\mathbf{x}_{1:i-1}]=\mathbf{A}_{i,:}\mathbf{x},$ (15)
$l_{i}^{2}$ :$\vcentcolon=\mbox{Var}[x_{i}|\mathbf{x}_{1:i-1}]=\mathbf{V}_{i,i}^{-2},$ (16)
@@ 4 Solving For Exponential Tilting Parameters
Botev (2017) proposed to find the optimal $\bm{\gamma}$, denoted by $\hat{\bm{\gamma}}$, by minimizing the worst likelihood ratio as shown in (10). Denoting $\log h$ by $\psi$, Botev (2017) showed that $\psi$ is convex with respect to (w.r.t.) $\bm{\gamma}$ and concave w.r.t. $\mathbf{y}$. Hence, finding $\hat{\bm{\gamma}}$ amounts to finding the solution $(\hat{\mathbf{y}},\hat{\bm{\gamma}})$ of the nonlinear system $\nabla\psi=\mathbf{0}$ if the solution satisfies $\mathbf{a}\leq\mathbf{L}\hat{\mathbf{y}}\leq\mathbf{b}$, which was empirically shown to be true by Botev (2017). Under our parameterization of $\psi$ that uses $\mathbf{x}$ instead of $\mathbf{y}$, it can be shown that the convex-concave property still holds.
@@
The integrand $h(\mathbf{x};\hat{\bm{\gamma}})=\exp\psi(\mathbf{x};\hat{\bm{\gamma}})$ has vanishing relative error:
@@
when the integration limits are $(\alpha\mathbf{a},+\bm{\infty})$ and $a_{i}>0$ for $i=1,\ldots,n$. Here, the variance is taken with respect to $\mathbf{x}$ sampled from $q(\mathbf{x};\hat{\bm{\gamma}})$.
The VRE property suggests that the proposal density $q(\mathbf{x};\hat{\bm{\gamma}})$ can become indistinguishable from the MVN density $\phi_{n}(\mathbf{x};\bm{\Sigma})$ truncated in the region of $(\alpha\mathbf{a},+\bm{\infty})$.
@@ 5.1 Variable Reordering Based on Vecchia
5:  Minimize (e.g., using L-BFGS) $g=\|\nabla\psi\|^{2}$ described in (18) with $\nabla g=\mathbf{H}\cdot\nabla\psi$ described in (18) and (19). Denote the solution by $(\hat{\mathbf{x}},\hat{\bm{\gamma}})^{\top}$
@@
In most settings, $\bm{\Sigma}$ is a kernel matrix that does not need to be stored explicitly and whose entries can be computed as needed based on the kernel and the relevant inputs or locations, resulting in a $\mathcal{O}(n)$ memory footprint. After variable reordering, VMET has a time complexity of $\mathcal{O}(Tnm+Nnm)$, where the summands correspond to optimizing the exponential tilting parameter $\bm{\gamma}$ and drawing MC samples of the VMET integrand, respectively. The number of iterations $T$ for optimizing $\bm{\gamma}$ typically ranges between $500$ and $1{,}000$, while the MC sample size $N$ is typically between $10^{4}$ and $10^{5}$. Although Vecchia-based variable reordering has quadratic complexity in $n$, it is usually not needed repeatedly for estimating the unknown parameters in $\bm{\Sigma}$ to achieve a smooth likelihood surface that is beneficial to optimization, as will be discussed in Section 7.1.
@@ 6 Numerical comparison
The original implementation of SOV in the R package mvtnorm (Genz & Bretz, 2009) only allows input MVN dimensions $n$ up to $1{,}000$; for $n>1{,}000$ we used the SOV implementation in tlrmvnmvt (Cao et al., 2022). The implementations used for MET, TLR, and VCDF were the TruncatedNormal (Botev, 2017) R package, the tlrmvnmvt (Cao et al., 2022) R package, and the R package from https://github.com/Recca2012/CDFApprox, respectively. tlrmvnmvt, TruncatedNormal, and our implemented VMET all utilized C/C++ source code through Rcpp (Eddelbuettel & François, 2011), while VCDF is based on a pre-compiled binary package. Computation times were measured on a standard scientific workstation.
@@ 6.1 Variable reordering comparison
Scenario 1 grid in $[0,1]^{2}$$\bm{\infty}$ $\mathbf{0}$
Scenario 2 Latin hypercube in $[0,1]^{2}$$\bm{\infty}$ $\mbox{Uniform}(-2,0)$
Scenario 3 grid in $[0,1]^{2}$ $-\mathbf{1}$ $\mathbf{1}$
Table 1: Spatial locations and integration limits for three simulation scenarios. $\mbox{Uniform}(-2,0)$ refers to generating random numbers from the uniform distribution in $(-2,0)$ independently for each coefficient.
@@ 6.2 Comparison for low-dimensional MVN probabilities
Overall, VMET had the best trade-off between computation cost and accuracy, achieving the same level of errors as MET while reducing the computation times by 90%. VCDF had the largest estimation errors, followed by TLR and SOV. TLR under-performed SOV because the TLR Cholesky factor of $\bm{\Sigma}$ introduced bias and the block variable reordering used in TLR is less effective in variance reduction than the univariate reordering. The third example had centered integration limits, where the optimal $\hat{\bm{\gamma}}$ for the proposal density used in MET and VMET was simply zero, hence the MET integrands were identical to the SOV integrands. For MET, solving for $\hat{\bm{\gamma}}$ dominated the computation cost, as seen from the computation time difference between MET and SOV, both at $N=10{,}000$, although the times were based on different implementations. Utilizing our proposed linear-complexity gradient and Hessian estimators in (18) and (19), the optimization cost of $\bm{\gamma}$ was significantly reduced, enhancing the scalability of the importance-sampling technique proposed in Botev (2017). We show in the next section that due to our complexity advantage, which is $\mathcal{O}(n)$ and $\mathcal{O}(n^{0.5})$ lower than SOV and TLR, respectively, our computation times were lower than those of SOV and TLR when $n$ was larger. The estimation bias of our method can be significant when $m$ is chosen too small (e.g., $m=10$ in Figure 2), but a reasonably large $m$ (e.g., between $30$ and $50$) was sufficient to achieve negligible bias in our comparisons, even in tens of thousands of dimensions.
@@ 6.3 Comparison for high-dimensional MVN probabilities
For larger $n$, we only compared our method to TLR and SOV, as MET and VCDF are computationally infeasible. We considered three $n=6{,}400$-dimensional MVN probabilities according to Scenarios 1 to 3 in Table 1, respectively. For this high-dimensional study, the nugget parameter was increased from $0.01$ to $0.03$ to avoid singularity, while the other covariance parameters remained unchanged. The RMSE of log-probability estimates and computation times are shown in Figure 3.
Figure 3: For the scenarios in Table 1 with $n=6{,}400$, we show the RMSE of 30 log-probability estimates against the computation time (both on a log scale), for VMET, TLR, and SOV. VMET is run with $m\in\{30,50,70\}$ and $N=10^{5}$, SOV and TLR are run with $N\in\{2,5,10\}\times 10^{4}$. For Scenarios 1 and 2, the average of VMET estimates at $m=70$ is used as truth. For Scenario 3, the average of SOV estimates at $N=10^{5}$ is used as truth.
@@ 6.5 Sampling from TMVN distributions
Our proposed VMET method can also be used for sampling TMVN distributions. We compared the performance of VMET with that of MET in terms of the quantile-quantile (q-q) plot of the drawn samples and the time used. For the TMVN distribution, we considered Scenario 1 in Table 1 with $n=900$ and the same Matérn kernel as in Sections 6.1 and 6.2. We generated $N=1{,}000$ samples using the VMET method with $m=30$ described in Algorithm 1 and the MET method implemented in the TruncatedNormal R package; a visual comparison is provided in Figure 5.
@@ 7 Analysis of Partially Censored Data
Notice that $\mathbf{x}_{c(i)}$ for $i>n_{1}$ may include observed $x_{j}$ (i.e., with $j\leq n_{1}$). We can view (21) as a limit case of (1) with $\mathbf{a}_{1:n_{1}}\rightarrow\mathbf{z}_{1:n_{1}}\leftarrow\mathbf{b}_{1:n_{1}}$ and $\mathbf{a}_{n_{1}+1:n}=-\bm{\infty}$.
@@ 7.1 Parameter estimation
It is common practice in many environmental applications to substitute censored data with a constant value, such as half the level of detection (LOD), the LOD divided by the square root of 2, or simply zero (Croghan & Egeghy, 2003). These LOD-based methods are also broadly used as a benchmark for more sophisticated Markov-chain-Monte-Carlo-based methods (De Oliveira, 2005; Ordoñez et al., 2018), which are often not scalable to large $n$. We show that in the presence of censored data, the censored MVN model in (21) can achieve more accurate parameter estimation than LOD-based methods. We generated a spatial field $\mathbf{x}$ over $6{,}400$ locations on a $80\times 80$ grid in the unit square based on a GP with mean zero and a Matérn covariance with variance, range, smoothness, and nugget of $(1.0,0.1,1.5,0.03)$ in the same setup as Scenario 1 in Table 1. All $x_{i}$ values below $b_{i}=0$ were censored to obtain the data $\mathbf{z}$. Because the censoring thresholds were zero, all LOD-based methods discussed in (Croghan & Egeghy, 2003) amounted to the same data augmentation.
@@
We considered a specific realization, for which $2{,}585$ responses were censored and $3{,}815$ were observed. To visualize the log-likelihood manifolds of the censored MVN model and the LOD approach, we fixed variance, smoothness, and nugget to their true values while plotting the log-likelihood against the range parameter. The GP log-likelihood using LOD-augmented data peaked at the lower bound of the interval we considered, whereas the censored MVN yielded a log-likelihood curve that peaked at exactly the true value $0.1$, indicating the efficacy gain from properly including the information at censored locations. For parameter estimation, we did not apply the variable reordering discussed in Section 5.1 to avoid discontinuity in the log-likelihood surface, because both the proposal density $q(\mathbf{x};\bm{\gamma})$ and the likelihood ratio $h(\mathbf{x};\bm{\gamma})$ are computed w.r.t. a given ordering of the integration variables.

### v4

# Linear-Cost Vecchia Approximation of Multivariate Normal Probabilities

@@ 1 Introduction
Current computational approaches for MVN probabilities are mostly based on the separation of variable (SOV) method (Genz, 1992). SOV applies a series of transformations such that the MVN probability can be written as an expectation of quantities with low variance, allowing an approximation of the expectation as the average of Monte Carlo (MC) samples. SOV consists of two stages: in the pre-processing stage, the Cholesky factor of the $n\times n$ MVN covariance matrix is computed in $\mathcal{O}(n^{3})$ time; in the MC sampling stage, $N$ samples are drawn, each in $\mathcal{O}(n^{2})$ time. The resulting SOV time complexity is $\mathcal{O}(n^{3}+Nn^{2})$. To improve upon the original SOV method, Gibson et al. (1994) introduced a univariate variable reordering technique used in the pre-processing stage that produces the Cholesky factor and further reduces the MC variance. Based on the series of integrand transformations used in the SOV method, Botev (2017) used importance sampling with minimax exponentially tilted (MET) normal proposal densities during the MC sampling stage to reduce variance. This results in substantially higher accuracy for tail MVN probabilities than SOV, but it requires iterative optimization of the proposal density in the pre-processing stage at a per-iteration complexity of $\mathcal{O}(n^{3})$; in practice, this optimization can be more expensive than the MC sampling for moderately large $n$ (e.g., n>1,000n>1{,}000). Sampling from truncated MVN (TMVN) distributions is closely connected to approximating MVN probabilities (Botev, 2017). Specifically, the integrand in MET (or SOV) can be decomposed into two parts that can be viewed as a proposal density and a density ratio, which can be used for importance sampling from TMVN distributions at the same cost as estimating the corresponding MVN probability. MET is currently the state-of-the-art in terms of achieving the highest accuracy for estimating MVN probabilities and the highest acceptance rate for sampling TMVN distributions.
@@ 2 Importance-Sampling Views of SOV and MET
$={}$ $\int_{\tilde{a}_{1}}^{\tilde{b}_{1}}\frac{\Phi(\tilde{b}_{1})-\Phi(\tilde{a}_{1})}{\Phi(\tilde{b}_{1})-\Phi(\tilde{a}_{1})}\phi(y_{1})\cdots\int_{\tilde{a}_{n}}^{\tilde{b}_{n}}\frac{\Phi(\tilde{b}_{n})-\Phi(\tilde{a}_{n})}{\Phi(\tilde{b}_{n})-\Phi(\tilde{a}_{n})}\phi(y_{n})\mathrm{d}\mathbf{y},$ (3)
$={}$ $\int_{\tilde{a}_{1}}^{\tilde{b}_{1}}\frac{\Phi(\tilde{b}_{1}-\gamma_{1})-\Phi(\tilde{a}_{1}-\gamma_{1})}{\Phi(\tilde{b}_{1}-\gamma_{1})-\Phi(\tilde{a}_{1}-\gamma_{1})}\frac{\phi(y_{1})}{\phi(y_{1}-\gamma_{1})}\phi(y_{1}-\gamma_{1})\cdots$ (4)
@@
$={}$ $E_{q^{B}(\mathbf{y};\bm{\gamma})}[h^{B}(\mathbf{y};\bm{\gamma})],\ \mbox{with}$ (6)
@@
where $\bm{\gamma}$ are the parameters for the proposal density used in MET, and $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function (PDF) and the cumulative probability function (CDF) of the univariate standard normal distribution, respectively. From the perspective of importance sampling, (3) and (5) correspond to the SOV and the MET methods, respectively. Specifically, $q^{B}(\mathbf{y};\bm{\gamma})$ is the proposal density and $h^{B}(\mathbf{y};\bm{\gamma})$ is the density ratio between $\phi_{n}(\mathbf{y};\mathbf{I}_{n})$ truncated in $\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}$ and $q^{B}(\mathbf{y};\bm{\gamma})$, up to a normalizing constant. MET samples $\mathbf{y}$ from $q^{B}(\mathbf{y};\bm{\gamma})$ and then approximates $E_{q^{B}(\mathbf{y};\bm{\gamma})}[h^{B}(\mathbf{y};\bm{\gamma})]$ by the sample average. Unlike TMVN distributions, $q^{B}(\mathbf{y};\bm{\gamma})$ can be sampled as a sequence of $n$ univariate truncated normal distributions as implied by (8). Hence, we can apply accept-reject sampling to TMVN distributions by using $q^{B}(\mathbf{y};\bm{\gamma})$ as the proposal density to draw samples from $\phi_{n}(\mathbf{x};\bm{\Sigma})$ truncated in $\mathbf{a},\mathbf{b}]$. SOV is a special case of MET when the exponential tilting parameters $\bm{\gamma}=\mathbf{0}$. Using
@@ 3.1 Conditional mean and variance under Vecchia
$\eta_{i}$ $\vcentcolon=E[x_{i}|\mathbf{x}_{1:i-1}]=\mathbf{A}_{i,:}\mathbf{x},$ (15)
$l_{i}^{2}$ $\vcentcolon=\mbox{Var}[x_{i}|\mathbf{x}_{1:i-1}]=\mathbf{V}_{i,i}^{-2},$ (16)
@@ 4 Solving For Exponential Tilting Parameters
[…]
@@
[…]
@@
[…]
@@ 5.1 Variable Reordering Based on Vecchia
[…]
@@
In most settings, $\bm{\Sigma}$ is a kernel matrix that does not need to be stored explicitly and whose entries can be computed as needed based on the kernel and the relevant inputs or locations, resulting in a $\mathcal{O}(n)$ memory footprint. After variable reordering, VMET has a time complexity of $\mathcal{O}(Tnm+Nnm)$, where the summands correspond to optimizing the exponential tilting parameter $\bm{\gamma}$ and drawing MC samples of the VMET integrand, respectively. The number of iterations $T$ for optimizing $\bm{\gamma}$ typically ranges between $500$ and 1,0001{,}000, while the MC sample size $N$ is typically between $10^{4}$ and $10^{5}$. Although Vecchia-based variable reordering has quadratic complexity in $n$, it is usually not needed repeatedly for estimating the unknown parameters in $\bm{\Sigma}$ to achieve a smooth likelihood surface that is beneficial to optimization, as will be discussed in Section 7.1.
@@ 6 Numerical comparison
The original implementation of SOV in the R package mvtnorm (Genz & Bretz, 2009) only allows input MVN dimensions $n$ up to 1,0001{,}000; for n>1,000n>1{,}000 we used the SOV implementation in tlrmvnmvt (Cao et al., 2022). The implementations used for MET, TLR, and VCDF were the TruncatedNormal (Botev, 2017) R package, the tlrmvnmvt (Cao et al., 2022) R package, and the R package from https://github.com/Recca2012/CDFApprox, respectively. tlrmvnmvt, TruncatedNormal, and our implemented VMET all utilized C/C++ source code through Rcpp (Eddelbuettel & François, 2011), while VCDF is based on a pre-compiled binary package. Computation times were measured on a standard scientific workstation.
@@ 6.1 Variable reordering comparison
Scenario 1 grid in $1]^{2} -\bm{\infty} 𝟎\mathbf{0}$
Scenario 2 Latin hypercube in $0,1]^{2} -\bm{\infty}$ $\mbox{Uniform}(-2,0)$
Scenario 3 grid in $0,1]^{2}$ $\mathbf{1}$ $\mathbf{1}$
Table 1: Spatial locations and integration limits for three simulation scenarios. $\mbox{Uniform}(-2,0)$ refers to generating random numbers from the uniform distribution in (2,0)(-2,0) independently for each coefficient.
@@ 6.2 Comparison for low-dimensional MVN probabilities
Overall, VMET had the best trade-off between computation cost and accuracy, achieving the same level of errors as MET while reducing the computation times by 90%. VCDF had the largest estimation errors, followed by TLR and SOV. TLR under-performed SOV because the TLR Cholesky factor of $\bm{\Sigma}$ introduced bias and the block variable reordering used in TLR is less effective in variance reduction than the univariate reordering. The third example had centered integration limits, where the optimal $\hat{\bm{\gamma}}$ for the proposal density used in MET and VMET was simply zero, hence the MET integrands were identical to the SOV integrands. For MET, solving for $\hat{\bm{\gamma}}$ dominated the computation cost, as seen from the computation time difference between MET and SOV, both at N=10, [^0] N=10{,}000, although the times were based on different implementations. Utilizing our proposed linear-complexity gradient and Hessian estimators in (18) and (19), the optimization cost of $\bm{\gamma}$ was significantly reduced, enhancing the scalability of the importance-sampling technique proposed in Botev (2017). We show in the next section that due to our complexity advantage, which is $\mathcal{O}(n)$ and $\mathcal{O}(n^{0.5})$ lower than SOV and TLR, respectively, our computation times were lower than those of SOV and TLR when $n$ was larger. The estimation bias of our method can be significant when $m$ is chosen too small (e.g., $m=10$ in Figure 2), but a reasonably large $m$ (e.g., between $30$ and $50$) was sufficient to achieve negligible bias in our comparisons, even in tens of thousands of dimensions.
@@ 6.3 Comparison for high-dimensional MVN probabilities
For larger $n$, we only compared our method to TLR and SOV, as MET and VCDF are computationally infeasible. We considered three n=6,400n=6{,}400-dimensional MVN probabilities according to Scenarios 1 to 3 in Table 1, respectively. For this high-dimensional study, the nugget parameter was increased from $0.01$ to $0.03$ to avoid singularity, while the other covariance parameters remained unchanged. The RMSE of log-probability estimates and computation times are shown in Figure 3.
Figure 3: For the scenarios in Table 1 with n=6,400n=6{,}400, we show the RMSE of 30 log-probability estimates against the computation time (both on a log scale), for VMET, TLR, and SOV. VMET is run with $m\in\{30,50,70\}$ and $N=10^{5}$, SOV and TLR are run with $N\in\{2,5,10\}\times 10^{4}$. For Scenarios 1 and 2, the average of VMET estimates at $m=70$ is used as truth. For Scenario 3, the average of SOV estimates at $N=10^{5}$ is used as truth.
@@ 6.5 Sampling from TMVN distributions
Our proposed VMET method can also be used for sampling TMVN distributions. We compared the performance of VMET with that of MET in terms of the quantile-quantile (q-q) plot of the drawn samples and the time used. For the TMVN distribution, we considered Scenario 1 in Table 1 with $n=900$ and the same Matérn kernel as in Sections 6.1 and 6.2. We generated N=1, [^0] N=1{,}000 samples using the VMET method with $m=30$ described in Algorithm 1 and the MET method implemented in the TruncatedNormal R package; a visual comparison is provided in Figure 5.
@@ 7 Analysis of Partially Censored Data
[…]
@@ 7.1 Parameter estimation
It is common practice in many environmental applications to substitute censored data with a constant value, such as half the level of detection (LOD), the LOD divided by the square root of 2, or simply zero (Croghan & Egeghy, 2003). These LOD-based methods are also broadly used as a benchmark for more sophisticated Markov-chain-Monte-Carlo-based methods (De Oliveira, 2005; Ordoñez et al., 2018), which are often not scalable to large $n$. We show that in the presence of censored data, the censored MVN model in (21) can achieve more accurate parameter estimation than LOD-based methods. We generated a spatial field $\mathbf{x}$ over 6,4006{,}400 locations on a $80\times 80$ grid in the unit square based on a GP with mean zero and a Matérn covariance with variance, range, smoothness, and nugget of $(1.0,0.1,1.5,0.03)$ in the same setup as Scenario 1 in Table 1. All $x_{i}$ values below $b_{i}=0$ were censored to obtain the data $\mathbf{z}$. Because the censoring thresholds were zero, all LOD-based methods discussed in (Croghan & Egeghy, 2003) amounted to the same data augmentation.
@@
We considered a specific realization, for which 2,5852{,}585 responses were censored and 3,8153{,}815 were observed. To visualize the log-likelihood manifolds of the censored MVN model and the LOD approach, we fixed variance, smoothness, and nugget to their true values while plotting the log-likelihood against the range parameter. The GP log-likelihood using LOD-augmented data peaked at the lower bound of the interval we considered, whereas the censored MVN yielded a log-likelihood curve that peaked at exactly the true value $0.1$, indicating the efficacy gain from properly including the information at censored locations. For parameter estimation, we did not apply the variable reordering discussed in Section 5.1 to avoid discontinuity in the log-likelihood surface, because both the proposal density $q(\mathbf{x};\bm{\gamma})$ and the likelihood ratio $h(\mathbf{x};\bm{\gamma})$ are computed w.r.t. a given ordering of the integration variables.